OpenStudy (anonymous):

If xy = -3 and (x + y)^2 = 33, what is the value of x^2 + y^2? (a) 27 (b) 30 (c) 33 (d) 36 (e) 39

OpenStudy (lgbasallote):

hint: (x+y)^2 = x^2 + y^2 + 2(xy)

OpenStudy (anonymous):

ok got the answer (a) 27

OpenStudy (lgbasallote):

how did you get that?

OpenStudy (anonymous):

i did x^2 + y^2 + 2(-3) = 33 which is x^2 + y^2 -6 =33 and then i soled for x^2 + y^2 which is 27. is that right?

OpenStudy (anonymous):

oh wait never mind it's 39 right?

OpenStudy (lgbasallote):

how did you get 27 if you had x^2 + y^2 - 6 = 33

OpenStudy (lgbasallote):

yup 39

OpenStudy (anonymous):

yeah i subtracted instead of adding but thanks